Klapperich Lab:Notebook/Lab Meeting Notes/2008/10/14

From OpenWetWare
Jump to: navigation, search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

14 October 2008 Lab Meeting Agenda

† MinCheol's research report today.

† Announcements

  • JZ and JD out this week at mTAS.

† Flu R01

  • TO DO More sample back? Can we put in 1.5 ml saline, collect in 3ml VTM tube directly? YES. Cathie will amend.
  • Plaque assay status?
  • TO DO Cathie STILL STILL has to submit the progress report to INSPIR.

† SEPSIS Project - CIMIT Bird Flu

  • B sub in blood is done. Lets talk about comepleting the gram positive plots so they "match."
  • SEM image analysis of pore size. DONE. Thanks JZ
  • Run E. coli RT-PCR using primers for a genomic target.
  • Report DNA amounts along with the RT-PCR data for the Sepsis paper.

†RNA project

  • Jeff Braman will visit in early Dec.

  • Do dilutions. Send to Bioanalyzer.


  • Mark finish control circuit
  • Fluorescence microscopy available from 720. Try beads experiment. DONE.
  • Bacteria Experiment?
  • Jane, Cathie, Larry, Ranjith met Friday. JZ will try process with modified washes.

† Valve Array/Valve building (FJ)

  • Waiting for fixture still from Melissa. Hopefully this week. UPDATE?
  • 9 Oct Shane says schematics are completed and would be emailed "right away". Have not received anything. Will contact 14 Oct to figure out why.

† Fraunhofer: LOAC

  • Update?
  • Alex working on more toluene bonding issues.
  • Try Ticona material. (HM, JD, AG)
  • Still Still Still troubleshooting ??

† Biointerfaces group

  • Mold still not made. Mask potentially Wed AM. When mask is made.
  • Do preliminary E. coli experiment soon. Sonali will train.

† CIMIT- Colson Grant

  • Preliminary plan for IRB submission at Harvard in place. Target date 1/1/09
  • STAR-CD Renewal - January ($2K)


  • QQ will run it on a gel to check.
  • Run Gels for water Wash steps - show me and Sonali UPDATE?
  • Only 1/5 chips are good at the hot embossing step. The vaccum chuck will be checked and o-rings changed - Will try after fixes and report next week. UPDATE?

* Fixuture for PCR chip. The heater and the thermocouples would be on the fixture.

  • Qingqing has to do experiments with MinCheol's designs. AS SOON AS CLEANING CONTEROLS ARE IN PLACE.


  • Cathie Still Still Working on Whitepaper.

† Senior project

  • Megan ordering oligos and primers for the RCA. Quote. Check with Heiko on his sources.
  • Haven't seen Teddy yet. What is up with him?

† F31: Cochlea

† Silica Optimization (Lambda):

  • Ongoing. More concentrations. .
  • Pressure testing on these channels when they are flowing.

hussam's update